### v1

# On Ruzsa’s conjecture on congruence preserving functions

On Ruzsa’s conjecture on congruence preserving functions
É. Delaygue Université Claude Bernard Lyon 1, CNRS, Centrale Lyon, INSA Lyon, Université Jean Monnet, ICJ UMR5208, 69622 Villeurbanne, France delaygue@math.univ-lyon1.fr
Résumé.
Ruzsa’s conjecture asserts that any sequence $(a_{n})_{n\geq 0}$ of integers that preserves congruences, i.e., satisfies $a_{n+k}\equiv a_{n}\mod k$, and has the growth condition $\limsup_{n\to+\infty}|a_{n}|^{1/n}<e$, must be a polynomial sequence. While previous results by Hall, Ruzsa, Perelli, and Zannier have confirmed this conjecture under stricter growth bounds, the general case remains open. In this paper, we establish a new partial result by proving that if in addition the generating series $f=\sum_{n\geq 0}a_{n}x^{n}$ has at most two singular directions at $x=0$, then $(a_{n})_{n\geq 0}$ is necessarily a polynomial sequence. Our approach is based on an adaptation of Carlson’s method, originally developed for the Pólya–Carlson dichotomy, combined with a refined analysis of Hankel determinants. Specifically, we derive an upper bound on these determinants using Pólya’s inequality and a transfinite diameter argument of Dubinin, while a non-Archimedean divisibility condition on Hankel determinants yields a lower bound, ultimately leading to the rationality of $f$. This confirms that counterexamples to Ruzsa’s conjecture, if they exist, must exhibit at least three singular directions.
Key words and phrases:
Congruence preserving functions, Pseudo-polynomials, Rusza’s conjecture, Polynomial sequences, Hankel determinants
1991 Mathematics Subject Classification:
11B50, 11C08 (Primary); 30B40, 11C20 (Secondary)
1. Introduction
Let $(a_{n})_{n\geq 0}$ be a sequence of integers that preserves congruences, that is, for every natural integers $n$ and $k$, we have
$a_{n+k}\equiv a_{n}\mod k.$
Such sequence is also known as a pseudo-polynomial [5] since, if $P\in\mathbb{Z}[x]$, then the sequence of integers $(P(n))_{n\geq 0}$ preserves congruences. We shall say that $(P(n))_{n\geq 0}$ is a polynomial sequence. A famous conjecture of Ruzsa says that there is a growth condition that forces pseudo-polynomials to be polynomial sequences.
Conjecture A (Ruzsa [9]).
Let $(a_{n})_{n\geq 0}$ be a pseudo-polynomial such that
$\underset{n\to+\infty}{\limsup}\,|a_{n}|^{1/n}<e.$
Then $(a_{n})_{n\geq 0}$ is a polynomial sequence.
This conjecture is sharp in the sense that Hall [5, p. 76] sketched an inductive construction of a pseudo-polynomial $(a_{n})_{n\geq 0}$ which is not a polynomial sequence but satisfies
$\underset{x\to+\infty}{\limsup}\,|a_{n}|^{1/n}\leq e.$
Ruzsa’s conjecture has been proven under more restrictive growth assumptions. In 1971, Hall [5] and Ruzsa [9] independently established this result when the bound $e$ is replaced by $e-1$. In 1984, Perelli and Zannier demonstrated that the conjecture holds if $e$ is replaced by $e^{0.66}$. Finally, in 1996, Zannier further improved this bound to $e^{0.75}$. To the best of our knowledge, no new cases of this conjecture have been proven since.
The results of Perelli and Zannier actually hold in a slightly more general setting, where congruences are considered only modulo prime numbers. Following [2], we define a sequence of integers $(a_{n})_{n\geq 0}$ as a primary pseudo-polynomial if, for every natural number $n$ and every prime $p$, we have
$a_{n+p}\equiv a_{n}\mod p.$
Perelli and Zannier’s approach first focused on proving the following elegant result.
Theorem A (Perelli–Zannier [7]).
Let $(a_{n})_{n\geq 0}$ be a primary pseudo-polynomial such that
$\underset{n\to+\infty}{\limsup}\,|a_{n}|^{1/n}<e.$
Then $(a_{n})_{n\geq 0}$ is $P$-recursive.
We recall that a sequence $(a_{n})_{n\geq 0}$ is said to be $P$-recursive if there exist $d\in\mathbb{Z}_{\geq 0}$ and polynomials $P_{0},\dots,P_{d}$ in $\mathbb{Q}[x]$, not all zero, such that, for every natural number $n$, we have
$P_{d}(n)\,a_{n+d}+\cdots+P_{0}(n)\,a_{n}=0.$
Equivalently [10], this means that the generating series $f$ of $(a_{n})_{n\geq 0}$ is annihilated by a nonzero linear differential operator with coefficients in $\mathbb{Q}[x]$, or, equivalently, that $f$ is $D$-finite.
Thus, under the assumptions of Theorem A, the generating series $f$ of $(a_{n})_{n\geq 0}$ is $D$-finite and has a strictly positive radius of convergence at the origin. In particular, it has only finitely many singularities in the complex plane, and we define the principal argument of any such singularity as a singular direction for $f$ at 00. The aim of this paper is to establish the following new partial case of Ruzsa’s conjecture.
Theorem 1.1.
Let $(a_{n})_{n\geq 0}$ be a primary pseudo-polynomial such that
$\underset{n\to+\infty}{\limsup}\,|a_{n}|^{1/n}<e.$
If its generating series has at most two singular directions at 00, then $(a_{n})_{n\geq 0}$ is a polynomial sequence.
Thus, a counterexample to Ruzsa’s conjecture could only exist if it has at least three singular directions at 00. Our approach is based on Carlson’s method [1] for proving the famous Pólya–Carlson dichotomy (see Theorem C below), a method that was recently used by Dimitrov [3] to prove the Schinzel–Zassenhaus conjecture on polynomials.
Zannier’s approach [11] to establishing Ruzsa’s conjecture with the bound $e^{0.75}$ instead of $e$ is of a different nature; it relies on deep results on the arithmetic of linear differential equations.
2. A Pólya–Carlson approach
Under the hypotheses of Theorem 1.1, we shall show that the generating series $f$ of $(a_{n})_{n\geq 0}$ is the power series expansion of a rational fraction. Using the congruences satisfied by $(a_{n})_{n\geq 0}$, it then becomes straightforward to prove that $f$ is the power series expansion of a rational fraction whose denominator is a power of $(1-x)$. This implies that $(a_{n})_{n\geq 0}$ is a polynomial sequence, as desired. In fact, Zannier showed in [11, p. 396] that if the generating series $f$ of a primary pseudo-polynomial $(a_{n})_{n\geq 0}$ is algebraic over $\mathbb{Q}(x)$, then $(a_{n})_{n\geq 0}$ is a polynomial sequence.
2.1. Kronecker’s rationality criterion
Let $f=\sum_{n=0}^{\infty}a_{n}x^{n}\in\mathbb{C}[[x]]$ be a formal power series with complex coefficients. For every positive integer $n$ we denote $H_{n}(f)$ the $n$-th Hankel matrix of $f$ defined by
$H_{n}(f)=\left(\begin{array}[]{cccc}a_{0}&a_{1}&\cdots&a_{n-1}\\ a_{1}&a_{2}&\cdots&a_{n}\\ \vdots&\vdots&&\vdots\\ a_{n-1}&a_{n}&\cdots&a_{2n-2}\end{array}\right).$
We shall say that $\det H_{n}(f)$ is the $n$-th Hankel determinant of $f$. Kronecker proved in [6] the following.
Theorem B (Kronecker’s rationality criterion).
A power series $f\in\mathbb{C}[[x]]$ is the expansion of a rational fraction if and only if almost all Hankel determinants of $f$ are zero.
When $f\in\mathbb{Z}[[x]]$, all Hankel determinants of $f$ belong to $\mathbb{Z}$, which provides non-archimedean bounds: for every nonnegative integer $n$ and every prime $p$, we have
$|\det H_{n}(f)|_{p}\leq 1,$
where $|\cdot|_{p}$ stands for the usual $p$-adic valuation on $\mathbb{Q}$. If one manages to prove that $|\det H_{n}(f)|<1$ for large enough $n$, then we have $\det H_{n}(f)=0$ for such $n$ and $f$ is the power series expansion of a rational fraction. This archimedean upper bound can be achieved via Pólya’s inequality under additional assumptions. This was Carlson’s idea [1] to prove the following dichotmoy.
Theorem C (Pólya–Carlson dichotomy).
A power series $f\in\mathbb{Z}[[x]]$ that converges inside the unit disk is either rational or admits the unit circle as a natural boundary.
Our approach to proving Theorem 1.1 is similar. We will first establish an archimedean upper bound on the Hankel determinants of $f$ using Pólya’s inequality. We will then derive a non-archimedean bound using the congruences satisfied by $(a_{n})_{n\geq 0}$, which will allow us to conclude.
2.2. An archimedean upper bound
The aim of this section is to prove the following archimedean upper bound.
Lemma 2.1.
Let $(a_{n})_{n\geq 0}$ be a sequence of complex numbers such that its generating series is $D$-finite with a positive radius of convergence $\rho$ and at most $r$ singular directions at 00. Then, we have
$\underset{n\rightarrow+\infty}{\limsup}\,|\det H_{n}(f)|^{1/n^{2}}\leq\frac{1}{4^{1/r}\rho}.$
The proof of the Pólya–Carlson dichotomy and of Lemma 2.1 is based on the following well-known inequality.
Theorem D (Pólya’s inequality [8]).
Let $(a_{n})_{n\geq 0}$ be a sequence of complex numbers and $f$ its generating series. Let $K$ be a compact subset of $\mathbb{C}$ such that $\mathbb{C}\setminus K$ is connected and
$g(x)=\sum_{n=0}^{\infty}a_{n}x^{-n-1}$
is analytic on $\mathbb{C}\setminus K$. Then, we have
$\underset{n\rightarrow+\infty}{\limsup}\,|\det H_{n}(f)|^{1/n^{2}}\leq\tau(K),$
where $\tau(K)$ is the transfinite diameter of $K$.
In our context, we follow the approach of Dimitrov [3] and will use the following theorem of Dubinin to estimate the transfinite diameter of a hedgehog
$\mathcal{K}(a_{1},\dots,a_{r})=\bigcup_{i=1}^{r}\left[0,a_{i}\right],$
where $a_{1},\dots,a_{r}$ are complex numbers.
Theorem E (Dubinin [4]).
The hedgehog $\mathcal{K}(a_{1},\dots,a_{r})\subset\mathbb{C}$ has transfinite diameter at most $(\max_{i}|a_{i}|^{r}/4)^{1/r}$, with equality if and only if the points $a_{1},\dots,a_{r}$ form the vertices of a regular $r$-gon centered at the origin.
Proof of Lemma 2.1.
Since $f$ is $D$-finite with a positive radius of convergence $\rho$ and at most $r$ singular directions, it can be analytically continued to the entire complex plane, except along a finite number of pairwise disjoint radial half-lines originating from some singularities $\alpha_{1},\dots,\alpha_{r}$ satisfying
$\min_{i}\,|\alpha_{i}|=\rho.$
Making the change of variables $x\mapsto 1/x$, we obtain that
$\frac{1}{x}f\left(\frac{1}{x}\right)=\sum_{n=0}^{\infty}a_{n}x^{-n-1}$
can be analytically continued to $\mathbb{C}\setminus\mathcal{K}$ where $\mathcal{K}$ is the hedgehog
$\mathcal{K}\left(\frac{1}{\alpha_{1}},\dots,\frac{1}{\alpha_{r}}\right)=\bigcup_{i=1}^{r}\left[0,\frac{1}{\alpha_{i}}\right].$
We can then use Theorems D and E to obtain the upper bound
$\underset{n\rightarrow+\infty}{\limsup}\,|\det H_{n}(f)|^{1/n^{2}}\leq\frac{1}{4^{1/r}\rho},$
as desired. ∎
2.3. A non-archimedean bound
The aim of this section is to prove the following non-archimedean bound.
Lemma 2.2.
Let $(a_{n})_{n\geq 0}$ be a primary pseudo-polynomial and $f$ its generating series. Then, for every positive integer, $\det H_{n}(f)$ is divisible by
$\prod_{\begin{subarray}{c}p\leq n-1\\ p\text{ prime}\end{subarray}}p^{n-p}.$
The main idea behind Lemma 2.2 is to consider the binomial transform $g$ of $f$, for which it is easier to obtain large divisibility properties of its Hankel determinants. If $(a_{n})_{n\geq 0}$ is a sequence of complex numbers, then we consider its binomial transform $(b_{n})_{n\geq 0}$ defined by
$b_{n}:=\sum_{k=0}^{n}(-1)^{n-k}\binom{n}{k}a_{k}.$
One can also recover $(a_{n})_{n\geq 0}$ from $(b_{n})_{n\geq 0}$ via the formula
$a_{n}=\sum_{k=0}^{n}\binom{n}{k}b_{k}.$
In the remainder of the article, if $f$ denotes the generating series of a sequence $(a_{n})_{n\geq 0}$ of complex numbers and $g$ represents the generating series of the binomial transform $(b_{n})_{n\geq 0}$ of $(a_{n})_{n\geq 0}$, then we shall refer to $g$ as the binomial transform of $f$. It is well-known that, for every positive integer, we have $\det H_{n}(f)=\det H_{n}(g)$. We nevertheless provide a proof of this result for the sake of completeness.
Lemma 2.3.
Let $g$ denote the binomial transform of a power series $f$. Then, for every positive integer $n$, we have $\det H_{n}(g)=\det H_{n}(f)$.
Démonstration.
In the following, if $A$ is a matrix, we denote by $[A]_{i,j}$ its entry at position $(i,j)$. For every positive integer $n$, we consider the lower triangular $n\times n$-matrix $L_{n}$ defined by
$[L_{n}]_{i,j}=\begin{cases}(-1)^{i-j}\binom{i-1}{j-1},&j\leq i,\\ 0,&j>i.\end{cases}.$
Let $n$ be a fixed positive integer and write $M_{n}:=L_{n}H_{n}(f)L_{n}^{\top}$. We shall prove that $M_{n}$ is equal to $H_{n}(g)$. For all $i$ and $j$, $1\leq i,j\leq n$, we have
$[M_{n}]_{i,j}$ $=\bigl{[}L_{n}H_{n}(f)L_{n}^{\top}\bigr{]}_{i,j}$
$=\sum_{r=1}^{n}\sum_{s=1}^{n}[L_{n}]_{i,r}[H_{n}(f)]_{r,s}[L_{n}^{\top}]_{s,j}$
$=\sum_{r=1}^{i}\sum_{s=1}^{j}[L_{n}]_{i,r}[H_{n}(f)]_{r,s}[L_{n}]_{j,s}$
$=\sum_{r=1}^{i}\sum_{s=1}^{j}(-1)^{i-r}\binom{i-1}{r-1}a_{r+s-2}(-1)^{j-s}\binom{j-1}{s-1}$
$=\sum_{r=0}^{i-1}\sum_{s=0}^{j-1}(-1)^{i+j-2-r-s}\binom{i-1}{r}\binom{j-1}{s}a_{r+s}.$ (2.1)
The identity $(1+X)^{i-1}(1+X)^{j-1}=(1+X)^{i+j-2}$ yields for every natural integer $m$ the classical convolution identity
[…]
This allows us to rewrite the double sum (2.1) as
$[M_{n}]_{i,j}=\sum_{m=0}^{i+j-2}(-1)^{i+j-2-m}\binom{i+j-2}{m}a_{m}=b_{i+j-2},$
that is $M_{n}=H_{n}(g)$, as desired. Since $\det L_{n}=1$, this identity proves the lemma. ∎
We now have the necessary tools to prove Lemma 2.2.
Proof of Lemma 2.2.
Let $(a_{n})_{n\geq 0}$ be a pseudo-polynomial, $(b_{n})_{n\geq 0}$ its binomial transform, and let $f$ and $g$ be their respective generating series. According to [2, Theorem 1], for every natural integer $n$, $b_{n}$ is divisible by the product $P_{n}$ of the primes less than or equal to $n$.
Since for every positive integer $n$, $P_{n}$ divides $b_{n}$, the $i$-th row of $H_{n}(g)$ can be factored by $P_{i-1}$. It follows that $\det H_{n}(g)$ is divisible by
$\prod_{k=1}^{n-1}P_{k}=\prod_{\begin{subarray}{c}p\leq n-1\\ p\text{ prime}\end{subarray}}p^{n-p},$
and so is $\det H_{n}(f)$ by Lemma 2.3. ∎
We can now prove our main result.
2.4. Proof of Theorem 1.1
Let $(a_{n})_{n\geq 0}$ be a primary pseudo-polynomial such that
$\underset{n\to+\infty}{\limsup}\,|a_{n}|^{1/n}<e,$
and let $f$ be its generating series. According to the work of Perelli and Zannier [7], $f$ is $D$-finite with a radius of convergence $\rho>1/e$. Since $f$ has at most two singular directions at 00, we can apply Lemma 2.1 and obtain that
$\underset{n\rightarrow+\infty}{\limsup}\,|\det H_{n}(f)|^{1/n^{2}}<\frac{e}{2}.$ (2.2)
Furthermore, according to Lemma 2.2, for all $n$ in $\mathbb{N}$, $\det H_{n}(f)$ is divisible by
$\prod_{\begin{subarray}{c}p\leq n-1\\ p\text{ prime}\end{subarray}}p^{n-p}.$ (2.3)
It therefore remains to determine an asymptotic estimate of the quantity (2.3). Taking the natural logarithm, we get
$\log\,\prod_{p\leq n-1}p^{n-p}$ $=\sum_{p\leq n-1}(n-p)\log p$
$=n\sum_{p\leq n-1}\log p-\sum_{p\leq n-1}p\log p.$
Let
$\theta(x)=\sum_{p\leq x}\log p$
be the first Chebyshev function. Then
$n\sum_{p\leq n-1}\log p=n\,\theta(n-1).$
To handle the term $\sum_{p\leq n-1}p\log p$, we use Abel transformation. Let $(c_{n})_{n\geq 0}$ be given by $c_{n}=\log n$ if $n$ is prime and $c_{n}=0$ otherwise. Then we have
$\sum_{p\leq n-1}p\log p$ $=\sum_{k=0}^{n-1}k\,c_{k}$
$=n\sum_{k=0}^{n-1}c_{k}-\sum_{k=0}^{n-1}\left(\sum_{i=0}^{k}c_{i}\right)(k+1-k)$
$=n\theta(n-1)-\sum_{k=0}^{n-1}\theta(k),$
leading to
$\log\,\prod_{p\leq n-1}p^{n-p}=\sum_{k=0}^{n-1}\theta(k).$
The Prime Number Theorem states that $\theta(x)\sim x$ as $x\to+\infty$. Hence, as $n\to+\infty$, we have
$\sum_{k=0}^{n-1}\theta(k)\sim\frac{n^{2}}{2},$
and
$\prod_{p\leq n-1}p^{n-p}=\exp\left(\frac{n^{2}}{2}+o(n^{2})\right).$ (2.4)
[…]
Aknowledgements. We warmly thank Dang-Khoa Nguyen for bringing Dubinin’s result to our attention.
Références
  • [1] F. Carlson, Über Potenzreihen mit ganzzahligen Koeffizienten, Math. Z. 9 (1921), no. 1-2, 1–13.
  • [2] É. Delaygue, T. Rivoal, On primary pseudo-polynomials (Around Ruzsa’s Conjecture), Int. J. Number Theory 18 (2022), no. 7, 1613–1636.
  • [3] V. Dimitrov, A proof of the Schinzel–Zassenhaus conjecture on polynomials, preprint, arXiv:1912.12545[math.NT] (2019).
  • [4] V. N. Dubinin, On the change in harmonic measure under symmetrization, Math. USSR, Sb. 52 (1985), 267–273.
  • [5] R. R. Hall, On pseudo-polynomials, Mathematika 18 (1971), 71–77.
  • [6] L. Kronecker, Zur Theorie der Elimination einer Variabeln aus zwei algebraischen Gleichungen, Berl. Monatsber. (1881), 535–600.
  • [7] A. Perelli, U. Zannier, On recurrent mod $p$ sequences, J. Reine Angew. Math. 348 (1984), 135–146.
  • [8] G. Pólya, Über gewisse notwendige Determinantenkriteterien für die Fortsetzbarkeit einer Potenzreihe, Math. Ann. 99 (1928), no. 1, 687–706.
  • [9] I. Z. Ruzsa, On congruence preserving functions, Mat. Lapok 22 (1971), 125–134.
  • [10] R. P. Stanley, Differentiably finite power series, European J. Combin 1 (1980), no. 2, 175–188.
  • [11] U. Zannier, On periodic mod $p$ sequences and $G$-functions (On a conjecture of Ruzsa), Manuscr. Math. 90 (1996), no. 3, 391–402.